OpenStudy (anonymous):

s^2+8s over s^2+4s-32 divided by s over s+4...im going nutty trying to do this one?

OpenStudy (anonymous):

=s(s+8)/ (s-4)(s+8) ÷ s/s+4 =s/s-4 ÷ s/s+4 =s/s-4 * s+4/s =s-4/s+4

OpenStudy (anonymous):

u factor.. then simplify..

OpenStudy (anonymous):

Thanx, it was upside down- but definitely confirmed my answer: s+4/s-4